Grade 11Magical Mathematics[Interesting Approach]

There are 'n' people and there are 'n2' marbles on the whole.No two of them collect the same number of marbles and each person collects at least one marble.What is the maximum number of marbles any one of them can collect.

please provide me the solution as soon as possible.

nikhil sadish
15 Years agoGrade 11

1 Answer

vikas askiitian expert
ApprovedApproved Tutor Answer15 Years ago

let first peson gets minimum marble = 1

second gets  minimum 2 marbles , third gets minimum 3 marbles & so on.......

 

this is an AP series whose first term is 1 & common difference is 1 ...

 

now total marbles distributed  to n-1 people = k

 

k = [(n-1)/2][2+(n-2)] = n(n-1)/2

 

now , last person gets all remaining marbles ..

 

remaining marbles = total - distributed

                          = n2 - (n(n-1)/2)

                          = n(n+1)/2

this is the maximum number of marble which any 1 of them can get ...

approve if u like my ans